# Doubly Robust Alignment for Large Language Models

### v1

@@ 7 Discussion
This work introduces a novel doubly robust preference optimization (DRPO) for LLM fine-tuning. Our approach enables accurate preference evaluation and policy optimization, providing robustness against misspecifications in both the reference policy and the preference model. We formally establish that our preference evaluation estimator is both doubly robust (Corollary 5.6) and semiparametrically efficient (Corollary 5.7) and demonstrate that our optimization procedure yields policies with a small regret (Theorem 5.8), and a lower suboptimality bound than DPO and PPO (Theorem 5.10). Our empirical results verify and reinforce the theoretical advantages, showing that DRPO achieves more robust and often superior performance to the baselines in preference optimization (Section 6.2). effectively address the high variance potentially arising from importance sampling ratios. An interesting direction for future work is to explore extensions of DRPO within the NLHF framework, by generalizing the estimator to handle arbitrary pairwise comparisons beyond the reference policy.
@@ A Technical Proof
A.2 Auxiliary lemma for proving Theorem 5.5
Before proceeding to the proof of Theorem 5.5, we first introduce an auxiliary lemma.
Lemma A.1.
Under Assumption 5.1, with $n$ independent data tuple $W_{i}=(X_{i},Y_{i}^{(1)},Y_{i}^{(2)},Z_{i}),i=1,\ldots n$, the efficient influence function [see e.g., Tsiatis, , 2006, for the detailed definition] for $p^{*}(\pi)$ is given by $\frac{1}{n}\sum_{i=1}^{n}\psi(X_{i},Y_{i}^{(1)},Y_{i}^{(2)},Z_{i};\pi,\pi_{\textup{ref}},g^{*})-p^{*}(\pi)$, with $\psi$ defined in equation (8).
Proof of Lemma A.1.
@@
The proof follows from standard techniques in semi-parametric statistic; see e.g., Chapters 2 & 3 in Bickel et al., [1998] and Theorem 3.5 in Tsiatis, [2006]. See also the proof of Theorem 1 in Kallus and Uehara, 2020a . Specifically:
@@
×$\qquad\qquad\times\pi_{\text{ref}}(y_{1}|x)\pi_{\text{ref}}(y_{2}|x)\frac{\partial}{\partial\gamma}f_{\gamma}(x)\bigg{\}}$
@@
×$\qquad\times\pi_{\text{ref}}(y_{1}|x)\pi_{\text{ref}}(y_{2}|x)\left(\frac{z}{g^{*}(y_{1},y_{2},x)}-\frac{1-z}{1-g^{*}(y_{1},y_{2},x)}\right)\frac{\partial}{\partial\eta}g_{\eta}(y_{1},y_{2},x)f_{X}(x)\Bigg{\}}$
@@
×$\qquad\times\pi_{\text{ref}}(y_{2}|x)f_{X}(x)\left(\frac{z}{g^{*}(y_{1},y_{2},x)}-\frac{1-z}{1-g^{*}(y_{1},y_{2},x)}\right)\frac{\partial}{\partial\eta}g_{\eta}(y_{1},y_{2},x).$
@@
By definition, the tangent space $\mathcal{T}_{\mathcal{M}}$ is the linear closure of the set of score functions of the all one-dimensional submodels regarding $\mathcal{M}$ that pass through true parameter; see Definition 2 in Kallus and Uehara, 2020a . Based on the likelihood function in equation (17), we can explicitly calculate the tangent space of the data generating process model $\mathcal{M}$. In fact, the tangent space $\mathcal{T}_{\mathcal{M}}$ is a product space, which can be represented as $\mathcal{T}_{f}\bigoplus\mathcal{T}_{\pi}\bigoplus\mathcal{T}_{g}$, with $\mathcal{T}_{f},\mathcal{T}_{\pi},\mathcal{T}_{g}$ being the sets of score functions of all one-dimensional submodels passing through the marginal distribution $f_{X}(x)$, conditional distribution $\pi_{\text{ref}}$ and preference probability $g^{*}$. Take the calculation of $\mathcal{T}_{f}$ as an example. Consider a one-dimensional submodel $\{f_{\varepsilon}(x)\}$, defined as
@@
$\mathcal{T}_{\pi}$ $=$ $\left\{q(y_{1},x)+q(y_{2},x):\mathbb{E}_{y\sim\pi_{\text{ref}}}[q(y,x)|X=x]=0,\sum_{y}\pi_{\text{ref}}(y|x)q^{2}(y,x)<\infty\right\},$
@@
$\psi_{1}(w)$ ::=:= $\left(\frac{\pi(y_{1}|x)}{\pi_{\text{ref}}(y_{1}|x)}-\frac{\pi(y_{2}|x)}{\pi_{\text{ref}}(y_{2}|x)}\right)(z-g^{*}(y_{1},y_{2},x))\pi_{\text{ref}}(y_{1}|x)\pi_{\text{ref}}(y_{2}|x)f_{X}(x)$
@@
$\psi_{2}(y_{1},y_{2},x)$ ::=:= $\mathbb{E}_{y^{*}\sim\pi}\left\{g(y^{*},y_{1},x)+g(y^{*},y_{2},x)\right\}-2\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\textup{ref}(\bullet|x)}\\ y^{*}\sim\pi(\bullet|x)\end{subarray}}\left\{g(y^{*},y,x)\right\},$
$\psi_{3}(x)$ ::=:= $2\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\textup{ref}(\bullet|x)}\\ y^{*}\sim\pi(\bullet|x)\end{subarray}}\left\{g(y^{*},y,x)\right\}-2p^{*}(\pi).$
@@
A.3 Proof of Theorem 5.5
Let $\mathbb{E}_{n}$ denote the empirical average over the $n$ tuples $(X,Y^{(1)},Y^{(2)},Z)$ in the dataset $\mathcal{D}$. Accordingly, our estimator for $p^{*}(\pi)$ can be represented by $\mathbb{E}_{n}[\psi(w;\pi,\widehat{\pi}_{\text{ref}},\widehat{g})]$.
@@
$\|\frac{\widehat{\pi}_{\text{ref}}}{\pi_{\text{ref}}}-1\|$ $=$ $\left(\mathbb{E}\left[\frac{\widehat{\pi}_{\text{ref}}(Y^{(1)}|X)}{\pi_{\text{ref}}(Y^{(1)}|X)}-1\right]^{2}\right)^{1/2}.$
@@
From Lemma A.1, we know that $\mathbb{E}_{n}\psi(w;\pi,\pi_{\text{ref}},g^{*})$ is an unbiased estimator for $p^{*}(\pi)$ with variance equal to SEB. Since both $\widehat{\pi}_{\text{ref}}$ and $\widehat{g}$ are obtained from external models independent of $\mathcal{D}$, analogous to the proof of Lemma 4.1, we know that the first term $\mathrm{I}$ and the second term $\mathrm{II}$ have zero means. The third term $\mathrm{III}$ is the bias term. Therefore, we obtain the following bias-variance decomposition for $\text{MSE}(\widehat{p}_{\text{DR}})$:
@@
Since $g^{*}$ is bounded by $1$, under the coverage assumption (Assumption 5.1), we obtain that
@@
This finishes the proof of Theorem 5.5.
A.4 Proofs of Corollaries 5.6 and 5.7
The proofs of Corollaries 5.6 and 5.7 follow directly from the assertion of Theorem 5.5.
A.5 Proof of Theorem 5.8
@@
Additionally, following the proof of Theorem 5.5, the bias of the proposed preference evaluation estimator can be upper bounded by
@@
It remains to upper bound the empirical process term $\mathbb{E}\sup_{\pi\in\Pi}|\widehat{p}_{\textup{DR}}(\pi)-\mathbb{E}\widehat{p}_{\textup{DR}}(\pi)|$. Toward that end, we employ Corollary 5.1 in Chernozhukov et al., [2014]. To invoke this corollary, notice that
@@
    According to Assumption 5.4, $\Pi$ is a policy class with VC dimension $v$. Under Assumption 5.1, it follows from Lemma A.6 in Chernozhukov et al., [2014] that the function class =$\mathcal{F}=\{\psi(\bullet,\pi,\widehat{\pi}_{\textup{ref}},\widehat{g})|\pi\in\Pi\}$ also has a VC dimension of $v$.
@@
Consequently, an application of Corollary 5.1 in Chernozhukov et al., [2014] yields that
$\mathbb{E}\sup_{\pi\in\Pi}|\widehat{p}_{\textup{DR}}(\pi)-\mathbb{E}[\widehat{p}_{\textup{DR}}(\pi)]|$ $=$ $O\left(\frac{1}{\sqrt{n}}\sqrt{\frac{v}{\epsilon^{2}}\log^{-1}\epsilon^{2}}+\frac{v}{n}\log^{-1}\epsilon^{2}\right)$
@@
Combining equations (25), (26) and (A.5), we obtain for any $\pi\in\Pi$ that
@@
This completes the proof of Theorem 5.8.
A.6 Proof of Corollary 5.9
The proof of Corollary 5.9 follows directly from the assertion of Theorem 5.8.
A.7 Proof of Theorem 5.10
Suboptimality gap for DRPO: If the BT assumption holds, we have $g^{*}(y_{1},y_{2},x)=\sigma(r^{*}(y_{1},x)-r^{*}(y_{2},x))$ where $\sigma(x)=1/(1+e^{-x})$ is the sigmoid function. Since the sigmoid function is monotonically increasing, under the realizability assumption, $\pi^{*}$ which maximizes $J(\pi)$ also maximizes $p^{*}(\pi)$. This follows from the classical results on the maximum rank correlation estimator that has been widely studied in the econometrics literature [see e.g., han1987non, sherman1993limiting]. Therefore,
$p^{*}(\pi^{*})-p^{*}(\widehat{\pi})$ $=$ $\mathbb{E}_{y^{*}\sim\pi^{*},\widetilde{y}\sim\widehat{\pi},y\sim\pi_{\textup{ref}}}\left\{g^{*}(y^{*},y,x)-g^{*}(\widetilde{y},y,x)\right\}$
@@
$=$ $\mathbb{E}_{y^{*}\sim\pi^{*},\widetilde{y}\sim\widehat{\pi}}\left\{\sigma^{\prime}(\xi)(r^{*}(y^{*},x)-r^{*}(\widetilde{y},x))\right\}$
$\geq$ $C_{0}(J(\pi^{*})-J(\widehat{\pi})),$
where $C_{0}$ is some positive constant and $\xi$ is some real number between $r^{*}(y^{*},x)-r^{*}(y,x)$ and $r^{*}(\widetilde{y},x)-r^{*}(y,x)$. Here, the second equality follows from mean value theorem. The last equality follows from the identity that $\sigma^{\prime}(x)=\sigma(x)(1-\sigma(x))$, which is bounded away from zeroo under Assumption 5.2 that the reward is bounded by some constant. Thus, we obtain $J(\pi^{*})-J(\widehat{\pi})=O(\textup{Reg}(\widehat{\pi}))$ and the suboptimality gap for DRPO follows directly from the assertion in Theorem 5.8.
@@
Using the fact that $l(\widetilde{\pi})\geq l(\pi^{*})$ and $l_{n}(\widehat{\pi})\geq l_{n}(\widetilde{\pi})$, we obtain the following upper bound:
$J(\pi^{*})-J(\widehat{\pi})$$\leq$ $\mathbb{E}\left\{[J(\pi^{*})-l(\pi^{*})]+[l(\widetilde{\pi})-l_{n}(\widetilde{\pi})]+[l_{n}(\widehat{\pi})-l(\widehat{\pi})]+[l(\widehat{\pi})-J(\widehat{\pi})]\right\}$ (27)
@@
Finally, under assumption 5.2, the function class =$\mathcal{F}=\left\{\sum_{y}\widehat{r}(y,X)\pi(y|X)\big{|}\pi\in\Pi\right\}$ is bounded by a constant. Using similar arguments to the proof of Theorem 5, we can employ Corollary 5.1 in Chernozhukov et al., [2014] to show that
@@
Suboptimality gap for DPO-based algorithm: We need some additional technical conditions to prove the suboptimality gap for DPO-based algorithms. Recall that when BT-model holds, there exists a one-on-one correspondence between the policy and reward model [Rafailov et al., , 2023]. We further assume
Assumption A.2 (Realizability).
The oracle reward $r^{*}$ lies in the bounded reward function class =$\mathcal{R}=\{\beta\log(\pi(y|x)/\pi_{\textrm{ref}}(y|x))+\beta Z(x):\pi\in\Pi\}$ induced by the policy class $\Pi$.
Assumption A.3 (Coverage).
@@
Assumption A.4 (Suboptimality gap for oracle reward).
@@
Notice that both the realizability and the coverage in Assumptions A.2 and A.3 differ from those in the main text. Specifically, Assumption A.2 imposes the realizability assumption on the oracle reward rather than the optimal policy whereas Assumption A.3 is stronger than that in the main text by requiring the denominators of the IS ratios to be strictly positive.
@@
$\mathbb{E}_{n}l(\widetilde{\pi})-\mathbb{E}_{n}l(\widehat{\pi})-\mathbb{E}l(\widetilde{\pi})+\mathbb{E}l(\widehat{\pi})$ (31)
$\leq$ $\mathbb{E}l(\widehat{\pi})-\mathbb{E}l(\widetilde{\pi})$
$\leq$ $\mathbb{E}l(\widehat{\pi})-\mathbb{E}l^{*}(\widetilde{\pi})+\mathbb{E}l^{*}(\widetilde{\pi})-\mathbb{E}l(\widetilde{\pi})$
@@
Moreover, according to Corollary 5.1 in Chernozhukov et al., [2014], using similar arguments to the proof of Theorem 5 and PPO-based algorithms, we have
$\mathbb{E}_{n}l(\widetilde{\pi})-\mathbb{E}_{n}l(\widehat{\pi})-\mathbb{E}l(\widetilde{\pi})+\mathbb{E}l(\widehat{\pi})$$\leq$ $2\mathbb{E}\sup_{\pi\in\Pi}|l(\pi)-\mathbb{E}l_{n}(\pi)|$ (32)
@@
where the last equality is due to that under Assumptions A.3 and A.4, the difference between $1$ and the ratio $\frac{\widehat{\pi}_{\textup{ref}}(y_{X}^{*}|X)\exp\left(\frac{1}{\beta}r^{*}(y_{X}^{*},X)\right)}{\sum_{y}\widehat{\pi}_{\textup{ref}}(y|X)\exp\left(\frac{1}{\beta}r^{*}(y,X)\right)}$ is of the order $O\left(\frac{1}{\epsilon}\exp\left(-\frac{\bar{c}}{\beta}\right)\right)$, almost surely.
@@
for a fixed $y_{0}$. Notice that the RHS corresponds to the mean squared error between $\widehat{r}$ and $r^{*}$, up to a baseline term that is independent of $Y^{(1)}$. Without loss of generality, we can assume this baseline term $r^{*}(y_{0},X)-\widehat{r}(y_{0},X)$ this equal to zero without affecting the validity of the proof. This is because the true reward can be redefined as $r^{*}(\bullet,X)-r^{*}(y_{0},X)$, since it is equivalent up to a function independent of the response. Similarly, the estimated optimal policy $\widehat{\pi}(\bullet|x)$ computed by DPO can be represented using the difference $\widehat{r}(\bullet,x)-\widehat{r}(y_{0},x)$, and we can replace $\widehat{r}$ in (35) using this difference. Consequently, we obtain that $\sigma^{2}\geq\epsilon\|\widehat{r}-r^{*}\|^{2}$ and hence

### v2

@@ 7 Discussion
This work introduces a novel doubly robust preference optimization (DRPO) for LLM fine-tuning. Our approach enables accurate preference evaluation and policy optimization, providing robustness against misspecifications in both the reference policy and the preference model. We formally establish that our preference evaluation estimator is both doubly robust (Corollary 3) and semiparametrically efficient (Corollary 4) and demonstrate that our optimization procedure yields policies with a small performance gap (Theorem 5), and a lower suboptimality bound than DPO and PPO (Theorem 7). Our empirical results reinforce the theoretical advantages, demonstrating DRPO’s robustness to reference-policy misspecification (Table 3; Figure 4, left) and preference-model misspecification (Table 3; Figure 4, right).
@@ A.1 Proof of Lemma 1
A.2 Auxiliary lemma for proving Theorem 2
Before proceeding to the proof of Theorem 2, we first introduce an auxiliary lemma.
Lemma 8.
Under Assumption 1, with $n$ independent data tuple $W_{i}=(X_{i},Y_{i}^{(1)},Y_{i}^{(2)},Z_{i}),i=1,\ldots n$, the efficient influence function [see e.g., 31, for the detailed definition] for $p^{*}(\pi)$ is given by $\frac{1}{n}\sum_{i=1}^{n}\psi(X_{i},Y_{i}^{(1)},Y_{i}^{(2)},Z_{i};\pi,\pi_{\textup{ref}},g^{*})-p^{*}(\pi)$, with $\psi$ defined in equation (8).
Proof of Lemma 8.
@@ A.2 Auxiliary lemma for proving Theorem 2
The proof follows from standard techniques in semi-parametric statistic; see e.g., Chapters 2 & 3 in Bickel et al. [160] and Theorem 3.5 in Tsiatis [31]. See also the proof of Theorem 1 in [124]. Specifically:
@@
$\qquad\qquad\times\pi_{\text{ref}}(y_{1}|x)\pi_{\text{ref}}(y_{2}|x)\frac{\partial}{\partial\gamma}f_{\gamma}(x)\bigg\}$
@@
$\qquad\times\pi_{\text{ref}}(y_{1}|x)\pi_{\text{ref}}(y_{2}|x)\left(\frac{z}{g^{*}(y_{1},y_{2},x)}-\frac{1-z}{1-g^{*}(y_{1},y_{2},x)}\right)\frac{\partial}{\partial\eta}g_{\eta}(y_{1},y_{2},x)f_{X}(x)\Bigg\}$
@@
$\qquad\times\pi_{\text{ref}}(y_{2}|x)f_{X}(x)\left(\frac{z}{g^{*}(y_{1},y_{2},x)}-\frac{1-z}{1-g^{*}(y_{1},y_{2},x)}\right)\frac{\partial}{\partial\eta}g_{\eta}(y_{1},y_{2},x).$
@@
By definition, the tangent space $\mathcal{T}_{\mathcal{M}}$ is the linear closure of the set of score functions of the all one-dimensional submodels regarding $\mathcal{M}$ that pass through true parameter; see Definition 2 in [124]. Based on the likelihood function in equation (17), we can explicitly calculate the tangent space of the data generating process model $\mathcal{M}$. In fact, the tangent space $\mathcal{T}_{\mathcal{M}}$ is a product space, which can be represented as $\mathcal{T}_{f}\bigoplus\mathcal{T}_{\pi}\bigoplus\mathcal{T}_{g}$, with $\mathcal{T}_{f},\mathcal{T}_{\pi},\mathcal{T}_{g}$ being the sets of score functions of all one-dimensional submodels passing through the marginal distribution $f_{X}(x)$, conditional distribution $\pi_{\text{ref}}$ and preference probability $g^{*}$. Take the calculation of $\mathcal{T}_{f}$ as an example. Consider a one-dimensional submodel $\{f_{\varepsilon}(x)\}$, defined as
@@
Meanwhile, consider a one-dimensional submodel
$\pi_{\varepsilon}(y|x)=\pi_{\text{ref}}(y|x)(1+\varepsilon q(y,x)),$
where $q(y,x)$ satisfies $\sum_{x}q^{2}(x,y)\pi_{\text{ref}}(y|x)<\infty$. Since we require $\pi_{\varepsilon}(y|x)$ be a valid conditional probability mass function, it must satisfy $\sum_{y}\pi_{\varepsilon}(y|x)=1$ for any $x$, which indicates $\mathbb{E}_{y\sim\pi_{\text{ref}}}q(y|x)=0$ for all $x$. Then the score function with respect to $\varepsilon$ is given by
$\frac{d}{d\varepsilon}\Big|_{\varepsilon=0}\log\pi_{\epsilon}(y_{1}|x)\pi_{\epsilon}(y_{2}|x)=q(y_{1},x)+q(y_{2},x).$
Therefore, the tangent space for the reference policy $\pi_{\text{ref}}$ can be represented as
$\mathcal{T}_{\pi}=\left\{q(y_{1},x)+q(y_{2},x):\mathbb{E}_{y\sim\pi_{\text{ref}}}[q(y,x)|X=x]=0,\sum_{y}\pi_{\text{ref}}(y|x)q^{2}(y,x)<\infty\right\}.$
@@
$\psi_{1}(w)$ $:=$ $\left(\frac{\pi(y_{1}|x)}{\pi_{\text{ref}}(y_{1}|x)}-\frac{\pi(y_{2}|x)}{\pi_{\text{ref}}(y_{2}|x)}\right)(z-g^{*}(y_{1},y_{2},x))\pi_{\text{ref}}(y_{1}|x)\pi_{\text{ref}}(y_{2}|x)f_{X}(x)$
@@
$\psi_{2}(y_{1},y_{2},x)$ $:=$ $\mathbb{E}_{y^{*}\sim\pi}\left\{g(y^{*},y_{1},x)+g(y^{*},y_{2},x)\right\}-2\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\textup{ref}(\bullet|x)}\\ y^{*}\sim\pi(\bullet|x)\end{subarray}}\left\{g(y^{*},y,x)\right\},$
$\psi_{3}(x)$ $:=$ $2\mathbb{E}_{\begin{subarray}{c}y\sim\pi_{\textup{ref}(\bullet|x)}\\ y^{*}\sim\pi(\bullet|x)\end{subarray}}\left\{g(y^{*},y,x)\right\}-2p^{*}(\pi).$
@@
A.3 Proof of Theorem 2
[…]
@@ A.3 Proof of Theorem 2
$\left\|\frac{\widehat{\pi}_{\text{ref}}}{\pi_{\text{ref}}}-1\right\|$ $=$ $\left(\mathbb{E}\left[\frac{\widehat{\pi}_{\text{ref}}(Y^{(1)}|X)}{\pi_{\text{ref}}(Y^{(1)}|X)}-1\right]^{2}\right)^{1/2}.$
@@
From Lemma 8, we know that $\mathbb{E}_{n}\psi(w;\pi,\pi_{\text{ref}},g^{*})$ is an unbiased estimator for $p^{*}(\pi)$ with variance equal to SEB. Since both $\widehat{\pi}_{\text{ref}}$ and $\widehat{g}$ are obtained from external models independent of $\mathcal{D}$, analogous to the proof of Lemma 1, we know that the first term $\mathrm{I}$ and the second term $\mathrm{II}$ have zero means. The third term $\mathrm{III}$ is the bias term. Therefore, we obtain the following bias-variance decomposition for $\text{MSE}(\widehat{p}_{\text{DR}})$:
@@
Since $g^{*}$ is bounded by $1$, under the coverage assumption (Assumption 1), we obtain that
@@
This finishes the proof of Theorem 2.
A.4 Proofs of Corollaries 3 and 4
The proofs of Corollaries 3 and 4 follow directly from the assertion of Theorem 2.
A.5 Proof of Theorem 5
@@
Additionally, following the proof of Theorem 2, the bias of the proposed preference evaluation estimator can be upper bounded by
@@
It remains to upper bound the empirical process term $\mathbb{E}\sup_{\pi\in\Pi}|\widehat{p}_{\textup{DR}}(\pi)-\mathbb{E}\widehat{p}_{\textup{DR}}(\pi)|$. Toward that end, we employ Corollary 5.1 in [144]. To invoke this corollary, notice that
@@
    According to Assumption 4, $\Pi$ is a policy class with VC dimension $v$. Under Assumption 1, it follows from Lemma A.6 in [144] that the function class $\mathcal{F}=\{\psi(\bullet,\pi,\widehat{\pi}_{\textup{ref}},\widehat{g})|\pi\in\Pi\}$ also has a VC dimension of $v$.
@@
Consequently, an application of Corollary 5.1 in [144] yields that
$\mathbb{E}\sup_{\pi\in\Pi}|\widehat{p}_{\textup{DR}}(\pi)-\mathbb{E}[\widehat{p}_{\textup{DR}}(\pi)]|$ $=$ $O\left(\frac{1}{\sqrt{n}}\sqrt{\frac{v}{\epsilon^{2}}\log^{-1}\epsilon^{2}}+\frac{v}{n\varepsilon}\log^{-1}\epsilon^{2}\right)$ (27)
@@
Combining equations (25), (26) and (27), we obtain for any $\pi\in\Pi$ that
@@
This completes the proof of Theorem 5.
A.6 Proof of Corollary 6
The proof of Corollary 6 follows directly from the assertion of Theorem 5. Before proving Theorem 7, we discuss the tightness of the suboptimality upper bounds derived therein for PPO, DPO, and DRPO. For each algorithm, its gap contains three components:
  • A bias term induced by KL-regularization, which is proportional to $\beta$.
  • A statistical complexity term of the form $\sqrt{v/n}$, which depends on the sample size $n$ and the complexity measure $v$ of the policy class.
  • A reward/reference policy estimation error term.
Since the first term can be made arbitrarily small by choosing a sufficiently small $\beta$, we discuss the tightness of the second and the third terms below. For the second term, our upper bounds for PPO and DPO match the lower bounds developed in [150], indicating their tightness, Finally, our theoretical investigation reveals that under certain settings – e.g., when the reward is linear and the prompt distribution is multivariate Gaussian – the suboptimality gap of PPO depends linearly on the estimation error of the regression coefficient, which itself is proportional to the reward model estimation error. Meanwhile, for DPO, when there is a constant gap between the specified and oracle reference policy, the algorithm suffers from a constant suboptimality gap that will not converge to zero. This demonstrates the tightness of the third term.
A.7 Proof of Theorem 7
Suboptimality gap for DRPO: If the BT assumption holds, we have $g^{*}(y_{1},y_{2},x)=\sigma(r^{*}(y_{1},x)-r^{*}(y_{2},x))$ where $\sigma(x)=1/(1+e^{-x})$ is the sigmoid function. Since the sigmoid function is monotonically increasing, under the realizability assumption, $\pi^{*}$ which maximizes $J(\pi)$ also maximizes $p^{*}(\pi)$. This follows from the classical results on the maximum rank correlation estimator that has been widely studied in the econometrics literature [see e.g., 161, 162]. Therefore,
[…]
@@ A.7 Proof of Theorem 7
[…]
$\geq$ $),\displaystyle C_{0}(J(\pi^{*})-J(\widehat{\pi}$)),
where $C_{0}$ is some positive constant and $\xi$ is some real number between $r^{*}(y^{*},x)-r^{*}(y,x)$ and $r^{*}(\widetilde{y},x)-r^{*}(y,x)$. Here, the second equality follows from mean value theorem. The last equality follows from the identity that $\sigma^{\prime}(x)=\sigma(x)(1-\sigma(x))$, which is bounded away from zeroo under Assumption 2 that the reward is bounded by some constant. Thus, we obtain $)J(\pi^{*})-J(\widehat{\pi})=O(\textup{Gap}(\widehat{\pi}$)) and the suboptimality gap for DRPO follows directly from the assertion in Theorem 5.
@@
[…]
@@
Finally, under assumption 2, the function class $\mathcal{F}=\left\{\sum_{y}\widehat{r}(y,X)\pi(y|X)\big|\pi\in\Pi\right\}$ is bounded by a constant. Using similar arguments to the proof of Theorem 5, we can employ Corollary 5.1 in [144] to show that
@@
Suboptimality gap for DPO-based algorithm: We need some additional technical conditions to prove the suboptimality gap for DPO-based algorithms. Recall that when BT-model holds, there exists a one-on-one correspondence between the policy and reward model [26]. We further assume
Assumption 5 (Realizability).
The oracle reward $r^{*}$ lies in the bounded reward function class $\mathcal{R}=\{\beta\log(\pi(y|x)/\pi_{\textrm{ref}}(y|x))+\beta Z(x):\pi\in\Pi\}$ induced by the policy class $\Pi$.
Assumption 6 (Coverage).
@@
Assumption 7 (Suboptimality gap for oracle reward).
@@
Notice that both the realizability and the coverage in Assumptions 5 and 6 differ from those in the main text. Specifically, Assumption 5 imposes the realizability assumption on the oracle reward rather than the optimal policy whereas Assumption 6 is stronger than that in the main text by requiring the denominators of the IS ratios to be strictly positive.
@@
[…]
@@
Moreover, according to Corollary 5.1 in [144], using similar arguments to the proof of Theorem 5 and PPO-based algorithms, we have
[…]
@@
where the last equality is due to that under Assumptions 6 and 7, the difference between $1$ and the ratio $\frac{\widehat{\pi}_{\textup{ref}}(y_{X}^{*}|X)\exp\left(\frac{1}{\beta}r^{*}(y_{X}^{*},X)\right)}{\sum_{y}\widehat{\pi}_{\textup{ref}}(y|X)\exp\left(\frac{1}{\beta}r^{*}(y,X)\right)}$ is of the order $O\left(\frac{1}{\epsilon}\exp\left(-\frac{\bar{c}}{\beta}\right)\right)$, almost surely.
@@
for a fixed $y_{0}$. Notice that the RHS corresponds to the mean squared error between $\widehat{r}$ and $r^{*}$, up to a baseline term that is independent of $Y^{(1)}$. Without loss of generality, we can assume this baseline term $r^{*}(y_{0},X)-\widehat{r}(y_{0},X)$ is equal to zero without affecting the validity of the proof. This is because the true reward can be redefined as $r^{*}(\bullet,X)-r^{*}(y_{0},X)$, since it is equivalent up to a function independent of the response. Similarly, the estimated optimal policy $\widehat{\pi}(\bullet|x)$ computed by DPO can be represented using the difference $\widehat{r}(\bullet,x)-\widehat{r}(y_{0},x)$, and we can replace $\widehat{r}$ in (36) using this difference. Consequently, we obtain that $\sigma^{2}\geq\epsilon\|\widehat{r}-r^{*}\|^{2}$ and hence